For a certain casino slot machine, the odds in favor of a win are given as 7 to 93. Express the indicated degree of likelihood as a probability value between 0 and1 inclusive The probability is Round to two decimal places as needed.)
Answer #1

Answer:

Odds of win 7 to 93

The odds in favor - the ratio of the number of ways that an outcome can occur compared to how many ways it cannot occur.

Odds in favor = Number of successes: 7

The odds against = 93

Probability of win = 7/(7+93) = 7/100

= 0.07